OpenStudy (anonymous):

write (2/3)^-3 in its simplest form

jhonyy9 (jhonyy9):

\[(2/3)^{-3} = 1/(2/3)^{3}\]

jhonyy9 (jhonyy9):

this is equal 27/8

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

yap!!! thank you so much!

jhonyy9 (jhonyy9):

than give me medal

jhonyy9 (jhonyy9):

thank you

OpenStudy (anonymous):

there you go

jhonyy9 (jhonyy9):

yw

OpenStudy (anonymous):

but.. wait i dont get why its not 8/27

OpenStudy (anonymous):

We solve this by using the following law of exponents ( p )^-n ( q )^n (---) = (---) ( q ) ( p ) Hence (2/3)^-3 = (3/2)^3 = (3)^3/(2)^3 = 27/8